# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ 4.3.2. Examples
Let $z\in\mathbb{R}$ be fixed and $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a linear operator. Consider the subspace $Y_{z}:=\mbox{span}\{(1,1,z)^{T}\}$. Let us first assume $z\neq 0$. We have that
$Y_{z}^{\perp}=\left\{(x,y,w)^{T}\in\mathbb{C}^{3}\mid w=\frac{-x-y}{z}\right\}.$
From the Definition 4.7 we have that the domain of $A_{Y_{z},\mathcal{L}}$ is made of functions $u\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=z\phi(L),\ z\left(\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right)\\ &\mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$ (4.64)
For instance, if we take $z=1$ and $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we have that $A_{Y_{1},\mathcal{L}}$ with domain
@@
$\begin{split}D(A_{Y_{1},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L)=\psi(L),\ \phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\ \psi^{\prime}(L)=0\mbox{ and }\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)=\alpha_{0}(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L))\}.\end{split}$ (4.65)
defines a contractile dynamic on the tadpole graph.
$-L$
$L$
$\psi$
$\phi$
Figure 3. Example of profile in (4.66)
On the other hand, assume $z=0$. In such case
$Y_{0}^{\perp}=\{(x,y,w)\in\mathbb{C}^{3}\mid x=-y\}.$
Therefore the domain of $A_{Y_{0},\mathcal{L}}$ is made of functions $\textbf{U}\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=0,\ \phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\ \mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$
In particular if we take $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we get that $A_{Y_{0},\mathcal{L}}$ with domain
$\begin{split}D(A_{Y_{0},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L),\ \psi(L)=0,\ \phi^{\prime}(L)=\phi^{\prime}(-L)\mbox{ and }\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\}.\end{split}$ (4.66)
defines an extension of $A_{0}$ generating a contractive dynamic on the tadpole.
@@ 5. The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped graphs
In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex. This strategy will allow us in the future to study the dynamics of NLS model (1.1) specially on $\mathcal{T}$-graphs.
@@ 5.1. Deficiency indices of the free Schrödinger operator
For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $-\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
@@
We note that $e^{\pm\frac{(1\pm i)}{\sqrt{2}}x}$ belong to $L^{2}(-L,L)$. On the other hand, on a half-line $[L,\infty)$, it would be square integrable only if the real part of ±$\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative. Along each edge $e=[L,\infty)$ we would therefore always have that the deficiency indices are ${d_{+}}_{e}={d_{-}}_{e}=1$.
@@ 5.2. Boundary systems
We define in the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by

### v4

@@ 4.3.3. Examples
Let $z\in\mathbb{R}$ and $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a contractive boundary relation. Consider the subspace $Y_{z}:=\operatorname{span}\{(1,1,z)^{\top}\}$.
Case $z\neq 0$$Y_{z}^{\perp}=\bigl\{(x,y,w)^{\top}\in\mathbb{C}^{3}:w=-(x+y)/z\bigr\}$. By Definition 4.9, the domain of $A_{Y_{z},R_{1}}$ consists of all $\mathbf{U}\in D(A_{0}^{*})$ satisfying
$\phi(-L)=\phi(L),\quad\psi(L)=z\,\phi(L),\quad z\!\left(\alpha_{1}\psi^{\prime\prime}(L)+\tfrac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\!\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right),$ (4.56)
together with $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$.
Taking $z=1$ and $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (which is contractive by Example 5), the domain specializes to
$\begin{split}D(A_{Y_{1},R_{1}})=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\phi(L)=\phi(-L)=\psi(L),\quad\phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\psi^{\prime}(L)=0,\quad\alpha_{1}\psi^{\prime\prime}(L)+\tfrac{\beta_{1}}{2}\psi(L)=\alpha_{0}\bigl(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\bigr)\Bigr\},\end{split}$ (4.57)
and by Theorem 4.10, $A_{Y_{1},R_{1}}$ generates a contraction semigroup on the tadpole graph.
[…]
$\phi(-L)=\phi(L),\quad\psi(L)=0,\quad\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L),$ (4.58)
together with $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$.
Taking $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, the domain specializes to
$\begin{split}D(A_{Y_{0},R_{1}})=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\phi(L)=\phi(-L),\quad\psi(L)=0,\\ &\phi^{\prime}(L)=\phi^{\prime}(-L),\quad\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\Bigr\},\end{split}$ (4.59)
and $A_{Y_{0},R_{1}}$ generates a contraction semigroup on the tadpole graph.
Example 7.
Assume $\alpha_{j}\equiv\alpha$ and $\beta_{j}\equiv\beta$ for all $j=0,\dots,N$. Let
$Y:=\operatorname{span}\{(0,0,1,\dots,1)^{T}\}\subset\mathbb{R}^{N+2},$
so that
$Y^{\perp}=\left\{(y_{1},y_{2},x_{1},\dots,x_{N})\in\mathbb{R}^{N+2}\;\Bigm|\;\sum_{j=1}^{N}x_{j}=0\right\}.$
For $N\in\mathbb{N}$ arbitrary and parameters $m_{1},\dots,m_{N}\in\mathbb{R}$ satisfying $m_{j}^{2}\leqslant N+1$ for all $j=1,\dots,N$, consider the operator-type boundary relation
$R_{1}:=\operatorname{graph}(\mathcal{L}),\qquad\mathcal{L}:=\operatorname{diag}(1,m_{1},\dots,m_{N}):\mathcal{K}_{+}\to\mathcal{K}_{-}.$ (4.60)
Its Krein adjoint relation is $R_{1}^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}=\frac{1}{N+1}\mathcal{L}$. The condition $m_{j}^{2}\leqslant N+1$ ensures that both $R_{1}$ and $R_{1}^{[*]}$ are $(\mathcal{K}_{+},\mathcal{K}_{-})$-contractions in the boundary system of Section 4.3, so that Theorem 4.10 applies.
The corresponding extension $A_{Y,R_{1}}$ is defined on
@@
$\begin{split}D(A_{Y,R_{1}}):=\Big\{\mathbf{U}\in H^{3}(\mathcal{G})\;\Bigm|\;&\phi(-L)=\phi(L)=0,\quad\psi_{1}(L)=\psi_{j}(L)\text{ for all }j=1,\dots,N,\\ &\phi^{\prime}(-L)=\phi^{\prime}(L)=m_{j}\psi_{j}^{\prime}(L)\text{ for all }j=1,\dots,N,\\ &\sum_{j=1}^{N}\psi_{j}^{\prime\prime}(L)=\frac{-\beta}{2\alpha}\,N\psi_{1}(L)\Big\},\end{split}$ (4.61)
and is the infinitesimal generator of a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$.
The special case $m_{j}=0$ for all $j=1,\dots,N$ corresponds to the boundary relation
$R_{1}^{0}:=\operatorname{graph}(\operatorname{diag}(1,0,\dots,0)),$
whose multivalued part satisfies $\operatorname{mul}(R_{1}^{0})=\{(0,y_{2},\dots,y_{N}):y_{j}\in\mathbb{C}\}$, reflecting the fact that the first-derivative traces $\psi_{1}^{\prime}(L),\dots,\psi_{N}^{\prime}(L)$ are forced to zero without any constraint on $\Lambda_{-}\mathbf{U}$. This is a genuine relation, not representable by any invertible operator, and the domain reduces to
$\begin{split}D(A_{Y,R_{1}^{0}})=\Big\{\mathbf{U}\in H^{3}(\mathcal{G})\;\Bigm|\;&\phi(L)=\phi(-L)=\phi^{\prime}(L)=\phi^{\prime}(-L)=0,\\ &\psi_{1}(L)=\psi_{j}(L)\text{ for all }j=1,\dots,N,\\ &\psi_{j}^{\prime}(L)=0\text{ for all }j=1,\dots,N,\\ &\sum_{j=1}^{N}\psi_{j}^{\prime\prime}(L)=\frac{-\beta}{2\alpha}\,N\psi_{1}(L)\Big\}.\end{split}$ (4.62)
The operator $A_{Y,R_{1}^{0}}$ generates a $C_{0}$-semigroup of contractions by Theorem 4.10, and its domain coincides with the looping-edge analogous of the class of vertex conditions employed in the study of stabilization and controllability of dispersive equations on star-shaped networks [6, 48].
4.4. A simple application to exponential stabilization
For simplicity we work within the setting of Example 6 (case $z=0$) on the tadpole graph $\mathcal{G}$ with $N=1$, taking $\alpha_{0}=\alpha_{1}=\alpha>0$ and $\beta_{0}=\beta_{1}=\beta\in\mathbb{R}$. The operator $A=A_{Y_{0},R_{1}}$ from (4.59), together with the relation $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, acts on all $\mathbf{U}=(\phi,\psi)\in H^{3}(\mathcal{G})$ satisfying
$\phi(-L)=\phi(L),\qquad\psi(L)=0,\qquad\phi^{\prime}(-L)=\phi^{\prime}(L),\qquad\phi^{\prime\prime}(-L)=\phi^{\prime\prime}(L),$
together with $\psi^{\prime}(L)=0$. This last condition follows directly from $R_{1}$: the equation $\Lambda_{-}\mathbf{U}=\mathcal{L}\,\Lambda_{+}\mathbf{U}$ reads
$\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}$,
and subtracting the two rows forces $\psi^{\prime}(L)=0$, after which either row gives $\phi^{\prime}(L)=\phi^{\prime}(-L)$. Fix $\gamma>0$ and consider the damped problem
$\partial_{t}\mathbf{U}(t)=A\,\mathbf{U}(t)-\gamma\,\mathbf{U}(t),\qquad t>0,\qquad\mathbf{U}(0)=\mathbf{U}_{0}\in D(A).$ (4.63)
Theorem 4.11.
Let $\mathbf{U}(t)$ be a classical solution of (4.63). Then the energy $E(t):=\tfrac{1}{2}\|\mathbf{U}(t)\|^{2}_{L^{2}(\mathcal{G})}$ satisfies
$E(t)\leqslant e^{-2\gamma t}\,E(0),\qquad t\geqslant 0,$
or equivalently $\|\mathbf{U}(t)\|_{L^{2}(\mathcal{G})}\leqslant e^{-\gamma t}\|\mathbf{U}_{0}\|_{L^{2}(\mathcal{G})}$ for all $t\geqslant 0$. In particular, the solution decays to zero exponentially at rate $\gamma$, independently of the dispersive coefficients $\alpha$ and $\beta$.
Proof.
We first address the existence of classical solutions to (4.63). Since $A_{Y_{0},R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ by Theorem 4.10, the operator $A-\gamma I$ is dissipative for every $\gamma>0$. Moreover, $(A-\gamma I)^{*}=A^{*}-\gamma I$, and since $A^{*}$ is also dissipative (as noted in Remark 4), so is $A^{*}-\gamma I$. The Lumer–Phillips theorem therefore guarantees that $A-\gamma I$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$, and in particular that (4.63) admits a unique classical solution $\mathbf{U}\in C^{1}([0,\infty);L^{2}(\mathcal{G}))\cap C([0,\infty);D(A))$ for every $\mathbf{U}_{0}\in D(A)$.
Writing (4.63) componentwise gives $\phi_{t}=-\alpha\phi^{\prime\prime\prime}-\beta\phi^{\prime}-\gamma\phi$ on (L,L)(-L,L) and $\psi_{t}=-\alpha\psi^{\prime\prime\prime}-\beta\psi^{\prime}-\gamma\psi$ on $(L,\infty)$. Differentiating the energy and substituting the equations of motion yields
$\frac{d}{dt}E(t)$ $=-\alpha\!\int_{-L}^{L}\!\phi\phi^{\prime\prime\prime}\,dx-\beta\!\int_{-L}^{L}\!\phi\phi^{\prime}\,dx-\gamma\!\int_{-L}^{L}\!\phi^{2}\,dx-\alpha\!\int_{L}^{\infty}\!\psi\psi^{\prime\prime\prime}\,dx-\beta\!\int_{L}^{\infty}\!\psi\psi^{\prime}\,dx-\gamma\!\int_{L}^{\infty}\!\psi^{2}\,dx.$
We claim that all terms except those with factor $\gamma$ vanish identically. Indeed, integrating by parts twice on the loop edge,
$\int_{-L}^{L}\phi\,\phi^{\prime\prime\prime}\,dx=\Bigl[\phi^{\prime\prime}\phi\Bigr]_{-L}^{L}-\int_{-L}^{L}\phi^{\prime\prime}\phi^{\prime}\,dx=\bigl(\phi^{\prime\prime}(L)\phi(L)-\phi^{\prime\prime}(-L)\phi(-L)\bigr)-\frac{1}{2}\bigl(\phi^{\prime}(L)^{2}-\phi^{\prime}(-L)^{2}\bigr),$
and every boundary term cancels upon applying $\phi(-L)=\phi(L)$, $\phi^{\prime}(-L)=\phi^{\prime}(L)$, and $\phi^{\prime\prime}(-L)=\phi^{\prime\prime}(L)$, giving $\int_{-L}^{L}\phi\,\phi^{\prime\prime\prime}\,dx=0$. Besides, since $\phi\phi^{\prime}=\tfrac{1}{2}(\phi^{2})^{\prime}$,
$\int_{-L}^{L}\phi\,\phi^{\prime}\,dx=\frac{1}{2}\bigl(\phi(L)^{2}-\phi(-L)^{2}\bigr)=0,$
again by $\phi(-L)=\phi(L)$. On the other hand, functions in $D(A)$ belong to $H^{3}(L,\infty)$, so $\psi$, $\psi^{\prime}$, and $\psi^{\prime\prime}$ all decay to zero as $x\to\infty$. Integrating by parts on the half-line therefore gives
$\int_{L}^{\infty}\psi\,\psi^{\prime\prime\prime}\,dx=\Bigl[\psi^{\prime\prime}\psi\Bigr]_{L}^{\infty}-\int_{L}^{\infty}\psi^{\prime\prime}\psi^{\prime}\,dx=-\psi^{\prime\prime}(L)\psi(L)+\frac{1}{2}\psi^{\prime}(L)^{2}=0,$
where the last equality uses $\psi(L)=0$ and $\psi^{\prime}(L)=0$. Similarly, $\int_{L}^{\infty}\psi\,\psi^{\prime}\,dx=\tfrac{1}{2}[\psi^{2}]_{L}^{\infty}=-\tfrac{1}{2}\psi(L)^{2}=0$. Consequently all boundary contributions vanish and the energy equation reduces to
$\frac{d}{dt}E(t)=-\gamma\|\mathbf{U}(t)\|^{2}_{L^{2}(\mathcal{G})}=-2\gamma\,E(t).$
An application of Grönwall’s lemma to the differential inequality $E^{\prime}(t)\leqslant-2\gamma E(t)$ gives the conclusion. ∎
Remark 5.
The vanishing of all boundary terms is not a coincidence. It is a direct manifestation of the fact that $A_{Y_{0},R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$: by Theorem 4.10, this is equivalent to the contractivity of $R_{1}$, which in turn forces the boundary terms arising in the Green identity to be non-positive. In the particular case at hand they vanish altogether, reflecting the precise balance between the subspace $Y_{0}$ and the relation $R_{1}$ in the definition of the domain. The systematic exploitation of this connection between the contractivity of the boundary relation, the structure of the domain, and the resulting energy dissipation is part of an ongoing work for the KdV equation on general metric graphs.
@@ 5. The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped graphs
In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering L-L as a terminal vertex. This strategy will allow us in the future to study the dynamics of NLS model (1.1) specially on $\mathcal{T}$-graphs.
@@ 5.1. Deficiency indices of the free Schrödinger operator
For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
@@
We note that $e^{\pm\frac{(1\pm i)}{\sqrt{2}}x}$ belong to $L^{2}(-L,L)$. On the other hand, on a half-line $L,\infty)$, it would be square integrable only if the real part of $\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative. Along each edge $e=[L,\infty)$ we would therefore always have that the deficiency indices are ${d_{+}}_{e}={d_{-}}_{e}=1$.
@@ 5.2. Boundary systems
We define on the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard skew-symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by
$\Lambda((\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),(\textbf{V},\mathcal{H}_{0}^{*}\textbf{V})):=[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}].$ (5.2)
Definition 5.2.
Let $\mathcal{J}=\mathbb{C}^{N+2}$. Define the standard skew-symmetric sesquilinear form on $\mathcal{J}\oplus\mathcal{J}$ by
$\rho\bigl((x_{1},y_{1}),(x_{2},y_{2})\bigr):=\langle y_{1},x_{2}\rangle_{\mathcal{J}}-\langle x_{1},y_{2}\rangle_{\mathcal{J}},\qquad(x_{i},y_{i})\in\mathcal{J}\oplus\mathcal{J},$ (5.3)
where $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathcal{J}}$ is the standard inner product on $\mathbb{C}^{N+2}$.
Definition 5.3.
We define $\mathbb{F}$ to be the surjective boundary map
$\begin{split}\mathbb{F}:&\ G(\mathcal{H}_{0}^{*})\to\mathcal{J}\oplus\mathcal{J}\\ &(\mathbf{U},\mathcal{H}_{0}^{*}\mathbf{U})\mapsto\bigl(\Gamma_{0}\mathbf{U},\,\Gamma_{1}\mathbf{U}\bigr),\end{split}$ (5.4)
where
$\Gamma_{0}\mathbf{U}=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\qquad\Gamma_{1}\mathbf{U}=\begin{pmatrix}\phi^{\prime}(-L)\\ -\phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}$.
Proposition 5.1.
The quintuple $(\Lambda,\mathcal{J},\mathcal{J},\mathbb{F},\rho)$ is a boundary system for $\mathcal{H}_{0}$, that is,
@@
Proof.
The proof follows by integration by parts. ∎
5.3. Self-adjoint extensions via boundary systems
The following theorem, taken from [49, Theorem 3.3], gives a complete parametrization of all self-adjoint extensions of $\mathcal{H}_{0}$ in terms of the boundary system constructed in the previous subsection.
Theorem 5.2.
An operator $\mathcal{H}$ is a self-adjoint extension of $\mathcal{H}_{0}$ if and only if there exist a subspace $X\subset\mathcal{J}=\mathbb{C}^{N+2}$ and a self-adjoint operator $\mathcal{L}:X\to X$ such that $\mathcal{H}=\mathcal{H}_{X,\mathcal{L}}$, where
$D(\mathcal{H}_{X,\mathcal{L}})=\Bigl\{\mathbf{U}\in H^{2}(\mathcal{G})\;\Big|\;\Gamma_{0}\mathbf{U}\in X,\quad\mathcal{L}\Gamma_{0}\mathbf{U}=P_{X}\Gamma_{1}\mathbf{U}\Bigr\}.$ (5.5)
Here $P_{X}:\mathcal{J}\to X$ is the orthogonal projection onto $X$.
Theorem 5.2 suggests a practical procedure for constructing self-adjoint extensions by prescribing boundary conditions in two stages. First, one translates the desired $k$ conditions (with $k<N+2$) into conditions on $\Gamma_{0}$ and $\Gamma_{1}$, which can then be directly incorporated into the definition of a matrix $M:X\to X$. The second stage consists of completing the remaining (unconstrained) entries of $M$ so that it becomes self-adjoint, and of determining the resulting conditions imposed by the relation $M\Gamma_{0}\mathbf{U}=P_{X}\Gamma_{1}\mathbf{U}$. We exemplify the procedure next.
5.4. Examples
Example 8 (Self-adjoint extensions on a looping-edge graph).
We seek all self-adjoint extensions of $\mathcal{H}_{0}$ whose domain includes only the condition looping condition $\phi(-L)=\phi(L)$. This condition constrains $\Gamma_{0}\mathbf{U}$ in a way that the condition $\Gamma_{0}\mathbf{U}\in X$ translates into the choice
$X:=\bigl\{x\in\mathcal{J}\;\big|\;x_{1}=x_{2}\bigr\},\qquad X^{\perp}=\operatorname{span}\{e_{1}-e_{2}\},\qquad\dim X=N+1,$
with orthogonal projection $P_{X}v=v-\tfrac{v_{1}-v_{2}}{2}(e_{1}-e_{2})$. Since no further constraint is imposed, $\mathcal{L}:X\to X$ is a free self-adjoint operator on $X$.
Consider the orthonormal basis for $X$
$g_{1}:=\frac{e_{1}+e_{2}}{\sqrt{2}},\qquad g_{k}:=e_{k+1},\quad k=2,\ldots,N+1.$
In this basis the coordinate vectors of $\Gamma_{0}\mathbf{U}$ and $P_{X}\Gamma_{1}\mathbf{U}$ are, respectively,
$\widetilde{\Gamma}_{0}\mathbf{U}=\bigl(\sqrt{2}\,\phi(-L),\;\psi_{1}(L),\;\ldots,\;\psi_{N}(L)\bigr)^{\!\top},\qquad\widetilde{\Gamma}_{1}\mathbf{U}=\Bigl(\tfrac{\phi^{\prime}(-L)-\phi^{\prime}(L)}{\sqrt{2}},\;\psi_{1}^{\prime}(L),\;\ldots,\;\psi_{N}^{\prime}(L)\Bigr)^{\!\top}.$
Write the matrix of $\mathcal{L}$ in the given basis in the following block form
$M=\begin{pmatrix}a&\mathbf{b}^{*}\\ \mathbf{b}&M^{\prime}\end{pmatrix},\qquad a\in\mathbb{R},\quad\mathbf{b}\in\mathbb{C}^{N},\quad M^{\prime}\in M_{N}(\mathbb{C})\text{ Hermitian},$ (5.6)
where $\mathbf{b}^{*}$ denotes the conjugate transpose of $\mathbf{b}$. This constitutes an $(N+1)^{2}$-real-parameter family.
Expanding $M\,\widetilde{\Gamma}_{0}\mathbf{U}=\widetilde{\Gamma}_{1}\mathbf{U}$ row by row one gets
$\phi^{\prime}(-L)-\phi^{\prime}(L)$ $=2a\,\phi(-L)+\sqrt{2}\,\mathbf{b}^{*}\vec{\psi}(L),$ (5.7)
$\vec{\psi}\,^{\prime}(L)$ $=\sqrt{2}\,\mathbf{b}\,\phi(-L)+M^{\prime}\vec{\psi}(L),$ (5.8)
where $\vec{\psi}(L)=(\psi_{1}(L),\ldots,\psi_{N}(L))^{\!\top}$ and $\vec{\psi}\,^{\prime}(L)=(\psi_{1}^{\prime}(L),\ldots,\psi_{N}^{\prime}(L))^{\!\top}$.
Theorem 5.3.
The self-adjoint extensions of $\mathcal{H}_{0}$ whose domain satisfies the looping condition $\phi(L)=\phi(-L)$ form an $(N+1)^{2}$-real-parameter family $\{\mathcal{H}_{a,\mathbf{b},M^{\prime}}\}$, parametrised by $a\in\mathbb{R}$, $\mathbf{b}\in\mathbb{C}^{N}$, and a Hermitian $M^{\prime}\in M_{N}(\mathbb{C})$ with domain
[…]
Remark 6.
The $2N+1$ real parameters $(a,\operatorname{Re}\mathbf{b},\operatorname{Im}\mathbf{b})$ account for two effects. The vector $\mathbf{b}\in\mathbb{C}^{N}$ couples the loop to the half-lines while the scalar $a\in\mathbb{R}$ introduces a ’Robin-type’ condition on the difference $\phi^{\prime}(-L)-\phi^{\prime}(L)$. Note in particular, setting $a=0$ and $\mathbf{b}=\mathbf{0}$ in (5.7)–(5.8) induces the pure periodic condition $\phi^{\prime}(-L)=\phi^{\prime}(L)$ together with $\vec{\psi}\,^{\prime}(L)=M^{\prime}\vec{\psi}(L)$ obtaining the fully decoupled self-adjoint extensions $(\mathcal{H}_{M^{\prime}},D(\mathcal{H}_{M^{\prime}}))$ with
$D(\mathcal{H}_{M^{\prime}})=\Bigl\{\mathbf{U}\in H^{2}(\mathcal{G})\;\Big|\;\phi(-L)=\phi(L),\quad\phi^{\prime}(-L)=\phi^{\prime}(L),\quad M^{\prime}\vec{\psi}(L)=\vec{\psi}\,^{\prime}(L)\Bigr\}.$ (5.9)
Several relevant extensions arise as special cases of $\mathcal{H}_{M^{\prime}}$:
  • $M^{\prime}=0$: the loop satisfies periodic boundary conditions and every half-line satisfies a Neumann condition $\psi_{j}^{\prime}(L)=0$ at the vertex.
  • $M^{\prime}=\alpha I_{N}$, $\alpha\in\mathbb{R}$: uniform Robin condition on the half-lines, $\psi_{j}^{\prime}(L)=\alpha\,\psi_{j}(L)$ for all $j$.
  • For a real non-zero parameter $Z$, the particular choice $M^{\prime}=\tfrac{1}{Z}\mathbf{1}\mathbf{1}^{T}$ (all entries equal to $\tfrac{1}{Z}$) is of particular interest for future works. Opening $\tfrac{1}{Z}\mathbf{1}\mathbf{1}^{T}\vec{\psi}(L)=\vec{\psi}^{\prime}(L)$ we get
    $\psi_{1}^{\prime}(L)=\cdots=\psi_{N}^{\prime}(L),\qquad\sum_{j=1}^{N}\psi_{j}(L)=Z\psi_{1}^{\prime}(L).$ (5.10)
    We obtain then the one-parameter family of decoupled self-adjoint extensions $\mathcal{H}_{Z}$ with domain
[…]
We now emulate the strategy of separating derivatives used for the Airy operator. We note in this case there is no actual separation of derivatives rather a simple reference frame to build some of the self-adjoint realizations.